### v3

# Optimal sensing policy with interference-model uncertainty

Abstract
Assume that an interferer behaves according to a parametric model but one does not know the value of the model parameters. Sensing enables to improve the model knowledge and therefore perform a better link adaptation. However, we consider a half-duplex scenario where, at each time slot, the communication system should decide between sensing and communication. We thus propose to investigate the optimal policy to maximize the expected sum rate given a finite-time communication. We first show that this problem can be modelled in the Markov decision process (MDP) framework. We then demonstrate that the optimal open-loop and closed-loop policies can be found significantly faster than with the standard backward-induction algorithm.
Index Terms:
Model uncertainty, sensing, MDP, link adaptation, superposition coding.
I Introduction
Dynamic sensing, spectrum access, and link adaptation have been widely studied in the literature. It consists in trying to predict the availability of the communication resource in order to optimize its usage. This approach is considered both in the scope of standard interference management and cognitive radio. In this latter case, the secondary user tries to avoid the primary user of a communication medium. While the applications of both fields are slightly different, the underlying technical approach is similar.
It generally involves some of the following steps: (Spectrum) sensing of the communication channel, interference prediction or primary-user behaviour prediction for cognitive radio cases, spectrum-access decision and link adaptation.
The optimal sensing duration, to optimize the rate, is for instance investigated in [8] in the scope of cognitive radio. Similarly, [12] investigates the optimal sensing time, but unlike [8] it uses an interference model. More specifically, in [12] sensing is performed at a low periodicity such that the rate is computed as a function of the expected number of missed transmissions or generated interference.
Regarding the interference prediction or primary-user detection, the Markov model approach is frequently used [13][6][2][9][12]. In general, it simply means that the interferer activation probability at time $t$ depends on its state at time $t-1$, where two states (active or inactive) are often considered.
We note that reinforcement learning (RL) was recently considered for this class of problems. For instance, [5] proposes a RL framework where an agent takes observations from the environment (e.g., from user equipments and the network) and learns proper policies that adjust the estimated SINR, such that a reward function (i.e., the user-equipment normalized throughput) is maximized. See also the survey [1] for more resources on this topic.
Limitations of existing works. Most of these studies consider that the parameters of the prediction models are perfectly known. They focus on sensing in order to optimize the usage of the interference model. They do not investigate the trade-off between interference-model improvement/learning via sensing and communication performance in a half-duplex case.
Considered problem. In this paper, similarly to some of the cited studies, we consider the sensing problem using a parametric interference model. However, we assume no apriori knowledge of the interference-model parameters.
We therefore aim to answer the following question: Given a communication duration of $T$ time slots, should one sense or communicate at any given time slot in order to maximize the expected sum rate? This corresponds for instance to a scenario with dynamic interference, where a new interferer is appearing every $T$ time slots.
We consider a simple model where an interferer is active with probability $p$ and not active with probability $1-p$ (in an independent and identically distributed (i.i.d) manner). The parameter $p$ is unknown at the beginning of the communication. The extension to the two-states Markov model is left for future work (see also the conclusions).
Structure of the paper and contributions. We first present the interference and channel models in Section II and characterize the model-parameter distribution. We then propose suited link-adaptation strategies, including superposition coding. In Section III, we first describe how the model can be presented as an MDP. We then describe efficient and optimal open-loop and closed-loop policies. Section IV provides simulation results to illustrate the performance of the proposed algorithm.
II Interference model, channel model, and link-adaptation strategies
As mentioned in the introduction, we consider the scenario where an interferer is active with probability $p$ and not active with probability $1-p$. The interferer draws a state according to $p$ and stays in the drawn state for a given time. Then, it switches to the newly sampled state. We assume that $p$ is not known and has to be estimated. More specifically, the system makes $n$ observations and observes $k$ times the channel with active interference. We assume an i.i.d. behaviour of this interferer.
II-A Interference characterization
Making these observations is equivalent to sampling $n$ times a Bernoulli random variable $Y$ with hidden parameter $p$. The corresponding binomial random variable is $Z=\sum_{i=1}^{n}Y_{i}$, and the probability to observe the interference channel $k$ times out of $n$ times for a given parameter $p$ is $P(Z=k)=\binom{n}{k}p^{k}(1-p)^{n-k}$. It follows that the random variable $p$ representing the parameter to estimate has therefore the distribution $p\sim f_{k,n}(x)$, with
[…]
This distribution is known as the Beta distribution [7], i.e., $p\sim Beta(k+1,n-k+1)$. Its expected value is $E[p]=\frac{k+1}{n+2}$.
In the paper, we use the notations $k^{\prime}=k+1$, $n^{\prime}=n+1$, $k^{\prime\prime}=k+2$, $n^{\prime\prime}=n+2$. We also have $p_{k}^{n}=\frac{k^{\prime}}{n^{\prime\prime}}$ and $\bar{p}_{k}^{n}=1-p_{k}^{n}$.
II-B Considered channel
We denote by: 1-$SINR_{1}$: The signal to interference and noise ratio (SINR) when the interferer is not active. 2-$SINR_{2}$: The SINR when the interferer is active.
We thus assume that the noise power in the considered additive white Gaussian noise (AWGN) channel is constant. Typically, $SINR_{2}<<SINR_{1}$. Thanks to Shannon theory, we know that the highest achievable rate on the AWGN channel (with any coding scheme) at a given SINR is $C(SINR)=1/2\log(1+SINR)$. Hence, the state of the interferer induces two channels. The first channel has a capacity $C(SINR_{1})=R_{max}$ and the second $C(SINR_{2})=R_{min}$. We also assume that the entire codeword experiences the same interference state.
II-C Link adaptation
Let $P$ be the average power per transmitted symbols, $N$ the power of the noise, and $I$ the power of the active interferer. Then, $N_{1}=N$ the power of the noise without active interference, and $N_{2}=N+I$ the noise and interference power, (i.e., case with active interference). Hence, $SINR_{i}=P/N_{i}$. Consider the following superposition-coding strategy, which uses two codes. It is similar to the one used for the broadcast channel [3][4, Chap. 15]. On the transmitter side, the sum of the two following codeword (from two distinct codes) is transmitted:
  • A first codeword with boosting (or power-allocation) constant $0\leq\alpha\leq 1$. The rate $R_{1}$ of the first code is $R_{1}(\alpha)=C(\frac{\alpha P}{N_{1}})$.
  • A second codeword with boosting constant $1-\alpha$. The rate $R_{2}$ of the second code is $R_{2}(\alpha)=C(\frac{(1-\alpha)P}{\alpha P+N_{2}})$
On the receiver side:
  • The receiver first tries to decode the second codeword. It is (almost) always successfully decoded, regardless of the interference.
  • The decoded codeword is subtracted to the received signal.
  • Then, the receiver tries to decode the first codeword: It is successfully decoded only in the case of no interference.
Hence, if there is no interference, the obtained rate is the sum of the rates of the two codes. If there is interference, the obtained rate is the rate of the second code.
Given the knowledge of $p$, the expected rate is therefore:
$\begin{split}E_{p}[R(\alpha,p)]=R_{1}(\alpha)(1-p)+R_{2}(\alpha).\end{split}$ (2)
Without the knowledge of $p$ but with $k,n$ observation values, the expected rate becomes:
[…]
Setting the derivative with respect to $\alpha$ to 0 yields the following optimal value
$\alpha_{k}^{n}=(N_{2}-N_{1}-N_{2}\cdot p_{k}^{n})/(P\cdot p_{k}^{n}).$ (4)
Then, $E_{f}[R(\alpha_{k}^{n},k,n)]$ is expected rate with the value of $\alpha$ that maximizes (3).
Note that the expected rate without superposition coding (using only one code per transmission) can be expressed as:
$\begin{split}E_{f}[R(\alpha_{k}^{n},k,n)]&=\max_{\alpha}\{E_{f}[R(\alpha=0)],E_{f}[R(\alpha=1)]\}.\\ &=\max\{R_{min},R_{max}\left(1-p_{k}^{n}\right)\}.\end{split}$ (5)
This can thus be seen as a constrained superposition-coding case, and therefore has inferior or equal performance. As an exemple, for $SINR_{1}=15$ dB, $SINR_{2}=10$ dB, and (known) $p=0.32$, the rate gain with superposition is around $12\ \%$.
III Optimal sensing strategy
As explained in the introduction, we consider a half-duplex model where the system should decide between sensing or communication at each time slot, to maximize the expected sum rate over $T$ time slots. We simplify the notation of (3) by using $R(\alpha_{k}^{n},k,n)$ for $E_{f}[R(\alpha_{k}^{n},k,n)]$.
III-A Modelization as an MDP
The system can be efficiently modelled as a finite-horizon Markov decision process (MDP) [10, Chap. 4]. The states, actions, rewards, and state-transition probabilities are the following.
  • States : The set $(k,n,i)$ defines a state $s(k,n,i)$ of the system, where $1\leq i\leq T$ is the slot index. The number of states is therefore:
    $\begin{split}&\sum_{i=1}^{T}(i\sum_{n=0}^{i}(n\sum_{k=0}^{n}k))=O(T^{6}).\end{split}$ (6)
  • Actions: Sensing $a=1$ or communication $a=0$.
  • (short-term) Rewards (which does not depend on $i$):
    $r(k,n,a=0)=R(\alpha_{k}^{n},k,n),\ r(k,n,a=1)=0.$ (7)
  • State-transition probabilities: For action $a=0$:
    $p(s(k,n,i+1)|s(k,n,i),a=0)=1,$ (8)
    and 0 for all other states. For action $a=1$:
    $\begin{split}&p(s(k+1,n+1,i+1)|s(k,n,i),a=1)=p_{k}^{n}\\ &p(s(k,n+1,i+1)|s(k,n,i),a=1)=\bar{p}_{k}^{n}.\end{split}$ (9)
As a result, the backward-induction algorithm can be used to optimally solve this MDP [10, Chap. 4]. However, the complexity is high due to the high number of states. We therefore propose an alternative approach to find the optimal strategy, whose complexity is significantly lower.
III-A1 Recalls on MDP and notations
The long-term gain is defined as $G_{i}=\sum_{j=i}^{T}r_{j}$, where $r_{j}$ is the short-term reward obtained at time $j$. The value of a state is defined by the expected gain starting from this state: $V(s(k,n,i))=E[G_{i}]$. Note that this is equal to the expected sum rate. For the sake of simplicity, we write the value of a state as $V(k,n,i)$ for $V(s(k,n,i))$. Under a given policy $\pi$, it is denoted as $V^{\pi}$ and under the optimal policy $\pi^{*}$ it is $V=V^{\pi^{*}}$. We use the notation $d_{i}=d(k,n,i)=a_{i}$ to denote the action in the sate $s(k,n,i)$ under the optimal policy $\pi^{*}=(d_{1},...d_{T})$. We shall also consider open-loop policies where a sequence of actions does not depend on the visited states, e.g., $\pi=(a_{1},...,a_{T})$.
III-A2 Similarities with bandit problems
The considered problem also has similarities with bandit problems [11, Chap.2], where the optimal exploitation-exploration trade-off should be found. When this framework is used to model a problem, unlike in the MDP framework, the reward does not in general depend on the state (see introduction of [11, Chap.3]).
The following greedy bandit algorithm ([11, Chap.2, Sec 2.2]), can be considered as benchmark: with probability $1-\epsilon$ take the optimal action with respect to the short-term reward and with probability $\epsilon$ do an unoptimal action (i.e., sensing in our case).
The proposed open-loop policy (see below) can also be seen as a bandit algorithm as the action does not depend on the state.
III-B Efficiently finding the optimal open and closed-loop policies
The two first subsections present intermediary results used to obtain the efficient open-loop policy (third subsection) and closed-loop policy (fourth subsection).
III-B1 Expected rate after $i$ sensing steps (open loop)
Assume that the current time is $i$ and $k$, $n$ are the current sensing values. $k_{i+1}$ denotes the number of interference observations at time $i+1$. The expected (instantaneous) rate after one additional sensing step is:
$\begin{split}E_{k_{i+1}}[R(k,n^{\prime})]=&p(k_{i+1}=k^{\prime}|k,n)R(\alpha_{k^{\prime}}^{n^{\prime}},k^{\prime},n^{\prime})+\\ &p(k_{i+1}=k|k,n)R(\alpha_{k}^{n^{\prime}},k,n^{\prime}),\end{split}$ (10)
where $p(k_{i+1}=k^{\prime}|k,n)=p_{k}^{n}$. The expected rate after $j$ additional sensing steps is:
$\begin{split}&E_{k_{i+j}}[R(k,n+j)]=g(k,n,j)=p_{k}^{n}g(k^{\prime},n^{\prime},j-1)+\bar{p}_{k}^{n}g(k,n^{\prime},j-1),\end{split}$ (11)
where $g(k,n,0)=R(\alpha_{k}^{n},k,n)$. Note that this corresponds to an open-loop mode meaning that sensing is performed regardless of the resulting states.
The following lemma shows that sensing always increases or keep constant the expected (instantaneous) rate.
Lemma 1.
For any $j\geq 0$, $g(k,n,j+1)\geq g(k,n,j)$.
Proof.
Take $j=0$. If no sensing is performed:
$\begin{split}&E_{k_{i+1}}[R(\alpha_{k}^{n},k,n)]=p_{k}^{n}R(\alpha_{k}^{n},k^{\prime},n^{\prime})+\bar{p}^{n}_{k}R(\alpha_{k}^{n},k,n^{\prime}).\end{split}$ (12)
If sensing is performed:
$E_{k_{i+1}}[R(\alpha_{k}^{n},k,n)]=p_{k}^{n}R(\alpha_{k^{\prime}}^{n^{\prime}},k^{\prime},n^{\prime})+\bar{p}_{k}^{n}R(\alpha_{k}^{n^{\prime}},k,n^{\prime}).$ (13)
where $R(\alpha_{k}^{n},k,n)\geq R(\alpha^{l}_{m},k,n)$ for any $l,m$ as, by definition, $\alpha_{k}^{n}$ is the best value given the $k$, $n$ sensing values. The rest of the proof uses the same argument for higher values of $j$. ∎
Computing (11) in a recursive manner yields an exponential complexity $O(2^{j})$. A better approach is the following. Equation (11) can be expressed as follows:
$\begin{split}&p_{k}^{n}g(k^{\prime},n^{\prime},j-1)+\bar{p}_{k}^{n}g(k,n^{\prime},j-1)\\ &=p_{k}^{n}(p_{k^{\prime}}^{n^{\prime}}g(k^{\prime\prime},n^{\prime\prime},j-2)+\bar{p}_{k^{\prime}}^{n^{\prime}}g(k^{\prime},n^{\prime\prime},j-2))\\ &+\bar{p}_{k}^{n}(p_{k}^{n^{\prime}}g(k^{\prime},n^{\prime\prime},j-2)+\bar{p}_{k}^{n^{\prime}}g(k,n^{\prime\prime},j-2))\\ &=p_{k}^{n}p_{k^{\prime}}^{n^{\prime}}g(k^{\prime\prime},n^{\prime\prime},j-2)+\\ &(p_{k}^{n}\bar{p}_{k^{\prime}}^{n^{\prime}}+\bar{p}_{k}^{n}p_{k}^{n^{\prime}})g(k^{\prime},n^{\prime\prime},j-2)+\bar{p}_{k}^{n}\bar{p}_{k}^{n^{\prime}}g(k,n^{\prime\prime},j-2).\end{split}$ (14)
where $g(k^{\prime},n^{\prime\prime},j-2)$ thus has to be computed only once instead of twice in the recursive approach. More generally, at level $j-m$ there are only $m+1$ distinct functions $g$ to evaluate after the merging step (corresponding to the last raw of (14)): indeed at this level the $k$-value (number of interference observations) is between $k$ and $k+m$. Algorithm 1 summarizes an efficient algorithm to compute $g(k,n,j)$. It has a complexity of $O(\sum_{l=0}^{j}l)=O(j^{2})$.
Algorithm 1 Computing the expected rate $E_{k_{i+j}}[R(k,n+j)]=g(k,n,j)$ (the expected rate after $j$ open-loop sensing steps)
1:  $p^{vec}$ = getProbaVec($k,n,j$)
2:  $g(k,n,j)=\sum_{k_{l}=k}^{k+j}p^{vec}_{k_{l}}R(\alpha_{k_{l}}^{n+j},k_{l},n+j)$
1:  Function $p^{vec}$ = getProbaVec(k,n,j)
2:  Initiate $p^{vec}$ as a size 1 vector of value 1.
3:  for $l=j:1$ do
[…]
5:     Get $p^{vec}$ of size $m+1$ corresponding to the probabilities of having $k_{l}=k,k_{l}=k+1,...,k_{l}=k+m$ (computed at step 7).
6:     Increment $n$ and for each $k_{l}$ value, multiply its corresponding probability by $p_{k_{l}}^{n}$ and $\bar{p}_{k_{l}}^{n}$, corresponding resp. to the values $k_{l}+1$ and $k_{l}$ (get a vector of probabilities of size $2(m+1)$).
7:     Merge the vector to obtain $p^{vec}$ with $m+2$ proba. values.
8:  end for
9:  Return $p^{vec}$.
III-B2 The on/off property of the problem
The following lemma shows that once sensing is stopped, one should always communicate. This is not surprising as the observation values $k$, $n$ do not change, but the remaining time to collect the gain (sum rate) diminishes.
Lemma 2.
Let $(a_{1},...,a_{T})$ be a realization of an optimal policy. If $a_{i}=0$, then $a_{j}=0$, for any $j>i$.
Proof.
We recall that Theorem 4.3.3 in [10] implies that the optimal policy is found by solving the optimality/Bellman equation (i.e., finding the state values satisfying the equation). The actions are then chosen according to this optimality equation, expressed as follows in our context:
$\begin{split}&V(k,n,i)=\max\{R(\alpha_{k}^{n},k,n)+V(k,n,i+1),\\ &p_{k}^{n}V(k^{\prime},n^{\prime},i+1)+\bar{p}_{k}^{n}V(k,n^{\prime},i+1)\}.\end{split}$ (15)
Assume that $a_{i-1}=d(k,n,i-1)=0$. It implies that
$\begin{split}&R(\alpha_{k}^{n},k,n)+V(k,n,i)>p_{k}^{n}V(k^{\prime},n^{\prime},i)+\bar{p}_{k}^{n}V(k,n^{\prime},i)\\ &\geq p_{k}^{n}Q(k^{\prime},n^{\prime},i,a=0)+\bar{p}_{k}^{n}Q(k,n^{\prime},i,a=0)\\ &=E_{k_{i}}[R(k,n^{\prime})]+p_{k}^{n}V(k^{\prime},n^{\prime},i+1)+\bar{p}_{k}^{n}V(k,n^{\prime},i+1),\end{split}$ (16)
where $Q(\cdot)$ is the standard $Q$-value and where $E_{k_{i}}[R(k,n^{\prime})]\geq R(\alpha_{k}^{n},k,n)$ as sensing always increase the expected short-term rate (Lemma 1). To develop the left-hand term, we assume that $a_{i}=d(k,n,i)=1$ (the opposite of the lemma for $j=i+1$):
$\begin{split}&R(\alpha_{k}^{n},k,n)+V(k,n,i)\\ &=R(\alpha_{k}^{n},k,n)+p_{k}^{n}V(k^{\prime},n^{\prime},i+1)+\bar{p}_{k}^{n}V(k,n^{\prime},i+1)\\ &>E_{k_{i}}[R(k,n^{\prime})]+p_{k}^{n}V(k^{\prime},n^{\prime},i+1)+\bar{p}_{k}^{n}V(k,n^{\prime},i+1),\\ \end{split}$ (17)
which yields $R(\alpha_{k}^{n},k,n)>E_{k_{i}}[R(k,n^{\prime})]$. We get a contradiction as $E_{k_{i}}[R(k,n^{\prime})]\geq R(\alpha_{k}^{n},k,n)$. Hence, the policy $(a_{1},..,a_{i-1}=0,a_{i}=0,a_{i+1},...)$ is better than $(a_{1},..,a_{i-1}=0,a_{i}=1,a_{i+1},...)$ given that $a_{i-1}=0$.
Then, using the same argument, we show that $(a_{1},...,a_{i-1}=0,a_{i}=0,a_{i+1}=0,...)$ is better than $(a_{1},..,,a_{i-1}=0,a_{i}=0,a_{i+1}=1,...)$. Applying successively this idea completes the proof of the lemma. ∎
III-B3 The optimal open-loop policy
This on/off property enables to get the following lemma. We recall that $g(k,n,j)$ can be efficiently computed with Algorithm 1.
Lemma 3.
Assume that the current time is $i-1$. Consider the following $(T-i)$ open-loop policies $\pi_{1}=(a_{i}=0,0,0,0,...)$, $\pi_{2}=(a_{i}=1,0,0,0,...)$, $\pi_{3}=(a_{i}=1,1,0,0,...)$, $\pi_{4}=(a_{i}=1,1,1,0,...)$, etc. Their corresponding expected gain is $g(k,n,0)(T-i)$, …, $g(k,n,j)(T-i-j)$, …, respectively.
If at least one $g(k,n,j)(T-i-j)>g(k,n,0)(T-i)=(T-i)R(\alpha_{k}^{n},k,n)$, with $j>0$, then the optimal action at time $i$ is sensing.
Note however that the lemma does not indicate the optimal action if $(T-i)R(\alpha_{k}^{n},k,n)$ has the largest value (i.e., it does not mean that communication is optimal).
Notice also that the above expected gains represent the values of the state $s(k,n,i)$ under the possible open-loop policies. Hence, the highest gain $g(k,n,j)(T-i-j)$ of the $j$-th open-loop policy can be used as an optimal open-loop policy, if one needs to set in advance the number of sensing time slots to be performed. As mentioned in Subsection III-A2, this open-loop policy can be seen as an efficient bandit policy.
Corollary 1.
The optimal open-loop policy is obtained by setting the number of sensing steps to $j$, $0\leq j\leq(T-i)$, such that $g(k,n,j)(T-i-j)$ has the highest value.
Proof.
(of Lemma 3) The equation of $V(k,n,i)$ is provided by (15). If $d(k,n,i)=0$, then the maximum is obtained with the first term of (15). Using Lemma 2, we get $Q(k,n,i,a=0)=(T-i)R(\alpha_{k}^{n},k,n)$. Hence, if we find at least one policy $\pi$ containing $d(k,n,i)=1$, which yields a value $V^{\pi}(k,n,i)>(T-i)R(\alpha_{k}^{n},k,n)$, then the optimal action is $d(k,n,i)=1$.
As a result, if for any $j\geq 0$ we have $g(k,n,j)(T-i-j)\geq(T-i)R(\alpha_{k}^{n},k,n)$, then
$V(k,n,i)\geq g(k,n,j)(T-i-j)\geq(T-i)R(\alpha_{k}^{n},k,n),$ (18)
where $V(k,n,i)$ is obtained with $d(k,n,i)=1$. ∎
III-B4 The optimal closed-loop policy
The value of the states under the optimal (closed-loop) policy can also be computed in reasonable time, using a slightly more advanced algorithm of similar complexity.
The main idea is to include a pre-processing part before the standard backward-induction algorithm such that this latter algorithm can be executed only on the subset of useful states.
Let $f(a,b,c)=\max\{a+b,c\}$. We re-express $V(k,n,i)$, given by (15) as
$\begin{split}&V(k,n,i)=f(p_{k}^{n}a_{i},\bar{p}_{k}^{n}b_{i},c_{i}),\end{split}$ (19)
with $a_{i}=V(k^{\prime},n^{\prime},i),\ b_{i}=V(k,n^{\prime},i),\ c_{i}=(T-i)R(\alpha_{k}^{n},k,n)$. The terms $a_{i}$ and $b_{i}$ are not known yet and therefore need to be further developed. We let $i^{\prime}=i+1$.
$\begin{split}&a_{i}=f(p_{k^{\prime}}^{n^{\prime}}a_{i^{\prime}},\bar{p}_{k^{\prime}}^{n^{\prime}}b_{i^{\prime}},c_{i^{\prime}}),\ b_{i}=f(p_{k}^{n^{\prime}}d_{i^{\prime}},\bar{p}_{k}^{n^{\prime}}e_{i^{\prime}},f_{i^{\prime}}).\end{split}$ (20)
Again, $c_{i^{\prime}}$ and $f_{i^{\prime}}$ can be directly computed using Lemma 2. The terms $a_{i^{\prime}}$, $b_{i^{\prime}}$, $d_{i^{\prime}}$, and $e_{i^{\prime}}$ need to be further expended. However, we have $b_{i^{\prime}}=d_{i^{\prime}}=V(k^{\prime},n^{\prime\prime},i+2)$, similarly to the observation made in (14). Hence, we have only three terms to expand: $a_{i^{\prime}},b_{i^{\prime}},e_{i^{\prime}}$. The connections between these two levels can be stored as a vector [1, 2, 2, 3]1223[1,\ 2,\ 2,\ 3][ 1 , 2 , 2 , 3 ], where the value at position $j$ indicates the index of the next-level state to be used. More generally, at level $i+j$ there are $j$ “fixed terms” to store and the connection vector has the form [1, 2, 2, 3, $[1,\ 2,\ 2,\ 3,\ 3,...,j+1]$.
Once level $i+j=T$ is reached, one can go in the other direction using the standard backward-induction algorithm starting from the obtained final states.
Algorithm 2 summarizes the two parts. The first part is used to identify only the useful states and useful connections.
Algorithm 2 Computing the value of $V(k,n,i)$ (under the optimal closed-loop policy)
1:  Start with the current $k$,$n$ values.
2:  First part (pre-processing)
3:  for $l=i:T$ do
[…]
5:     Store the $j+1$ fixed terms, each computed as $(T-l)R(\alpha_{k}^{n},k,n)$.
6:     Store the connections between this level and the next one.
7:  end for
8:  Second part (backward induction)
9:  for $l=T:i$ do
[…]
11:     Compute the value of the $j$ states at this level using:
  • The state values and the connection vector at the previous level $i+j+1$.
  • The fixed terms computed in the first part.
12:  end for
Regarding the complexity, similarly to Algorithm 1, there are are $O(j)$ operations to do per level. As a result, the complexity is $O(\sum_{l=i}^{T}l)=O(T^{2}-i^{2})$.
IV Simulation results
Figure 1 shows the values of a state $s(k=0,n=0,i=0)$ under several policies, with and without superposition coding:
  • The open-loop policies: For the $j$-th open-loop policy, the value is computed as $g(k,n,j)(T-i-j)$. Recall that Corollary 1 indicates that the highest value determines the optimal number of open-loop sensing steps.
  • The optimal closed-loop policy (computed with Algorithm 2), presented for several maximum allowed number of sensing steps.
  • The greedy (bandit) benchmark algorithm, with $\epsilon^{*}=0.01$.
Note that value of the state $s(k=0,n=0,i=0)$ gives the expected sum rate at the end of the communication. Hence, it directly gives the average sum rate that would be obtained when simulating the system for many interference values.
The value of the simulation parameters are the following: $T=1000$, $SINR_{1}=10$ dB, $SINR_{2}=15$ dB.
[…]
Figure 1: Value of the state $s(k=0,n=0,i=0)$ under several policies, $T=1000$. For the closed-loop policy the x-axis (number of sensing steps) represents the number of maximum allowed sensing steps.
Greedy ($\epsilon^{*}=0.03$) Open-loop Closed-loop
1 1.041 1.07
Figure 2: Normalized expected sum rate with $p\in[0.05,0.3]$, $T=500$, no superposition coding.
Finally, Figure 3 shows an example of evolution of a system over time, with $T=1000$. The first subplot shows the expected gain of each visited state (i.e., their value) over time, the second the chosen action, the third the current sum rate, and the last one the current interference-state realization. We observe that the algorithm (closed-loop) decided to switch from sensing to communication at time $i=27$, and the coding rate $R_{max}$ was chosen (example without superposition coding). Note that for more extreme values of $p$ (close to 0 or 1), yielding polarized $k$ with a high probability, the sensing period is significantly reduced, typically to less than 5 time slots. One possible interpretation is that a higher reliability on the estimate, for balanced values of $p$, is required due to a higher rate loss in case of a bad link-adaptation choice.
Figure 3: Example of evolution of a system over time, with $T=1000$. Only the first 60 slots are shown.
V Conclusions
In this paper, we presented a dilemma where the communication system should decide between communication or sensing to estimate the interference-model parameter. The latter operation enables to more accurately set the link adaptation, but at the cost of a missing communication opportunity. We modeled this problem as an MDP and showed that optimal policies, both open and closed loop, can be computed with a low computational complexity.
The optimal closed-loop policy is well approximated by the optimal open-loop policy.
Future work. If one replaces the considered interference model by a 2-states (interferer active and not active) Markov prediction model, which enables to perform predictions, then the states is described by $(k_{0},n_{0},k_{1},n_{1},\Delta,\alpha,i)$ where $k_{j},n_{j}$ characterize the transition from one state to the other state, $\Delta$ is the elapsed time since the last sensing operation, and $\alpha=\{0,1\}$ is the interference state at the last sensing operation. $\Delta$ and $\alpha$ (in addition to the model parameters) are used to compute the expected interference probability. With this framework, sensing is used not only to improve the model knowledge but also as an input to the prediction model. As a result, the on/off property (Lemma 2) does not hold anymore. Further research is therefore required to find a low-complexity optimal algorithm. Reinforcement learning may be considered.